4. Consider the following financial statement information for the Route 66 Company: Item Inventory Accounts receivable Accounts payable Beginning $1,273 3,782 1,795 Ending 1,401 3,368 2,025 Net sales Cost of goods sold $14,750 11,375 Calculate the operating and cash cycles.
Answer #1

Inventory turnover = $11,375/[(1,273 + 1,401) / 2] = 8.51 times

Receivables turnover = $14,750/[(3,782 + 3,368) / 2] = 4.13 times

Payables turnover = $11,375/[(1,795 + 2,025) / 2] = 5.96 times

Inventory period = 365 days / 8.51 times = 42.89 days

Receivables period = 365 days / 4.13 times = 88.38 days

Payables period = 365 days / 5.96 times = 61.24 days

So the time it takes to acquire inventory and sell it is about 43 days. Collection takes another 88 days, and the operating cycle is thus 43 + 88 = 131 days. The cash cycle is this 131 days less the payables period, 131 - 61 = 70 days.